Calculate the enthalpy change for this reaction given the following data. C(s) + 2 H2(g) → CH4(g)
ΔcH C(s) = -393, H2(g) = -286, CH4(g) = -890 kJ/mol
1
Expert's answer
2020-05-08T14:04:48-0400

Enthalpy change "= -393-2*286+890=890-393-572=-75 kJ"